Question 3 10.2.9-T
Test the hypothesis using the P value approach. Be sure the verify the requirements of the test.
H0:p=.46 versus H1: p<.46
n= 150 x=63 α=.01
I'm having a really hard time figuring out what p-value is

Answer by Boreal(15235)

This is a one sample proportion test. If all the assumptions are correct, the critical value is |z|>2.576
The proportion is 63/150, which is 0.42
z=(0.42-0.46)/sqrt (.46*.54)/150; sqrt term=0.0407
z=-0.04/0.0407=-0.9828
From the table or a calculator the probability left of that z-value is 0.16, which is the p-value. This is a one-way test, so it is not doubled. One would not reject Ho
I usually check these on the calculator, where a one sample proportion test is very easy to run.